Solid sodium sulfate, Na2SO4, is added slowly to a solution that is 1.0 x 10-4 in both Ba2+ and Pb2+ until [SO42-] reaches 1.0 x 10^-4 M. Would either BaSO4 (Ksp = 1.1 x 10^-10) or PbSO4 (Ksp = 1.7 x 10^-8) precipitate under these conditions?
Answer #1

[Ba^2+] = 1*10^-4M

[Pb^2+] = 1*10^-4M

BaSO4(aq) --------------> Ba^2+(aq) + SO4^2-(aq)

Ksp = [Ba^2+][SO4^2-]

1.1*10^-10 = [Ba^2+]*1*10^-4

[Ba^2+] = 1.1*10^-10/(1*10^-4) = 1.1*10^-6M

PbSO4(aq) --------------> Pb^2+(aq) + SO4^2-(aq)

Ksp = [Pb^2+][SO4^2-]

1.7*10^-8 = [Pb^2+]*1*10^-4

[Pb^2+] = 1.7*10^-8/(1*10^-4) = 1.7*10^-4M

Qsp of BaSO4 > Qsp of PbSO4

Ba^2+ first form precipitate